OpenStudy (anonymous):

sorry.. but what are two steps to solve these two problems? 2(x + 3) = 12 x/4 + 3 = 5 i am really bad at math..

OpenStudy (anonymous):

To solve: \[2(x + 3) = 12\] The first step is to FOIL that 2 by multiplying it by the x and 6

OpenStudy (anonymous):

You do that to get rid of the parentheses So it becomes: \[2x + 6 = 12\] Now all you have to do is solve for x which is 3

OpenStudy (deepolisnoob):

To make your equations simpler, you can divide the first equation by 2, eliminating the 2 on the first section and leaving you with 6 on the second section:\[\frac{ 2(x+3)=12 }{ 2 }\] You will be left with \[x+3=6\] Subtract 3 from both sides in order to have your x value on its own, leaving you with \[x=3\] ======================================================== In \[\frac{ x }{ 4 }+3=5\], you should get rid of the 3 first to keep only your fraction by itself, leaving you with \[\frac{ x }{ 4 }=5-3=2\] To get rid of your division sign, you multiply both sides by 4, giving you \[x=8\]

OpenStudy (anonymous):

The second problem: \[\frac{ x }{ 4 } + 3 = 5\] The first step is to multiply both sides of the equation by four: \[4 \times \frac{ x }{ 4 } + 3 = 5 \times 4\] We do that so we can get rid of that fraction and I problem becomes: \[x + 12 = 20\] Now solve for x which is 8 @sydneysalyer Was that great help or what would you like to become my fan to give you help when needed???